Find the unit normal to the surface 𝑦 = 𝑥 + 𝑧 3 at the point (1,2,1). 


1
Expert's answer
2021-11-23T16:54:43-0500
"F(x, y, z)=x-y+z^3=0"

"F_x=1, F_y=-1, F_z=3z^2"

Point "(1, 2, 1)"


"F_x=1, F_y=-1, F_z=3(1)^2=3"

"\nabla F=i-j+3k"

"|\nabla F|=\sqrt{(1)^2+(-1)^2+(3)^2}=\sqrt{11}"

"n=\dfrac{\nabla F}{|\nabla F|}=\dfrac{1}{\sqrt{11}}i-\dfrac{1}{\sqrt{11}}j+\dfrac{3}{\sqrt{11}}k"